OpenStudy (anonymous):

Solve the equation in the complex number system. x^3-1=0

OpenStudy (anonymous):

I got it down to (x-1)(x^2+1x+1), but I don't know how to finish it.

OpenStudy (campbell_st):

ok... so thats great you recongnised the difference of 2 cubes now use the general quadratic formula for the quadratic factor \[x^2 + x + 1\] means a = 1, b = 1 and c=1 substitute them into \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] you'll have a negative value in the square root... and use \[i^2 = -1 \] to manage it and then you'll be able to find the complex solutions. hope it helps

OpenStudy (anonymous):

Oh, duh! That makes perfect sense. Thank you. :)